OpenStudy (anonymous):

Find to the nearest ten thousandth, the radian mesasure of theta of a first quadrant angle with the given funtion value..... csc theta=2.346

OpenStudy (anonymous):

Do i just go onto my calculator and do the inverse of sin^-1(2.346)?

OpenStudy (anonymous):

1/sin^1. csc is the reciprocal of sin.

OpenStudy (anonymous):

so it would be sin^-1(1/sin2.346) and press enter right?

OpenStudy (anonymous):

@kropot72 Am I right?

OpenStudy (anonymous):

If you are familiar with series you can expand it as a taylor series and find the value with the required accuracy. Your way does work too

OpenStudy (anonymous):

Im not familiar with this"taylor series" but thank you for the clarity, I think ill use the calculator way, thank you!

OpenStudy (kropot72):

@Life What value did you get for theta?

OpenStudy (anonymous):

i didnt get it from anywhere its given

OpenStudy (kropot72):

The given value is the cosec of theta. What value for the angle theta did you get when you used your calculator as described?

OpenStudy (anonymous):

i actually got an error and so i did sin^-1(1/2.346) and i got a decimal from that, is it correct and my calculator is in radian mode

OpenStudy (kropot72):

You calculator must be in degree mode. You should get the value of 25.23028246 degrees for theta. To convert to radians divide the value of theta in degrees by 180 and multiply the result by pi.

OpenStudy (anonymous):

did you get .4403515...

OpenStudy (anonymous):

would i multiply by 180 and divide by pi for every question, like if it were regular trig funtions like sin cos and tan

OpenStudy (kropot72):

You are correct. Theta = 0.4403515 radians You must first find the value of the angle in degrees. This would require using the inverse trig function to find the angle in degrees if you have the sine, cos or tan etc. Next degrees are converted to radians by dividing by 180 and then multiplying the result by pi.

OpenStudy (anonymous):

what about if the question was cos theta=0.5379

OpenStudy (anonymous):

would i do exactly as you said above?

OpenStudy (kropot72):

\[\theta=\cos^{-1}0.5379\] You need to set your calculator to degree mode. The use the cos^-1 function to find theta. Please post the result here.

OpenStudy (anonymous):

57.459203...

OpenStudy (anonymous):

why am i using inverse if thats just a regular trig funtion?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Your result 57.459203 degrees is correct. You are given that cos theta=0.5379. Cosine is a trig ratio. To find the measure of theta in radians, given the cosine of theta, it is first necessary to find the value of theta in degrees. The meaning of the inverse trig function cos^-1 0.5379 is 'the angle in degrees that has the cosine value 0.5379'.

OpenStudy (anonymous):

So to conclude, with a general statement, when wanting to find out the radian measure in degrees, we always use the inverse trig func

OpenStudy (kropot72):

It is only necessary to use the inverse trig function when you are given the trig ratio of the angle concerned. For example if we are given sin theta = 1. Find the value of theta in radians: \[\sin^{-1} 1=90\ degrees\] \[\theta=90\ degrees=\frac{90}{180}\times \pi=\frac{\pi}{2}\ radians\]

OpenStudy (kropot72):

If it is given that theta = 180 degrees. Find theta in radians. \[\theta=180\ degrees=\frac{180}{180}\times \pi=\pi\ radians\]

OpenStudy (anonymous):

I got confused, please give me some time to understand everything.

OpenStudy (anonymous):

Back to the question where i asked for the cos of theta=0.5379, you said that the answer was 57.459203...but this is the answer before i divide by 180 and multiply by pi, isnt the final answer 1.00285?

OpenStudy (kropot72):

\[\frac{57.459203}{180}\times \pi=can\ you\ calculate?\]

OpenStudy (anonymous):

when you calculate the above, you get .3192177953, after that I multiply by pi and I get 1.002852281, is that the final answer because before you said that 57.459203 was the answer and I think you may have made a mistake.

OpenStudy (kropot72):

Yes the final answer is\[\theta=1.002852278\ radians\]

OpenStudy (anonymous):

Our answers are off

OpenStudy (anonymous):

the final three numbers are different, I got 281 and you got 278

OpenStudy (kropot72):

I confirmed that 57.459203 is the value of theta in degrees. I did not say that 57.459203 is the final answer. Do not worry about the difference in the three least significant figures of our results.

OpenStudy (kropot72):

The answer needs to be given to the nearest ten thousandth. In that case both our answers are; theta = 1.00285 radians

OpenStudy (anonymous):

My mistake, I just have one more question, how do I find the exact function value of an angle that is given in radians. For example cos 2(pi)/3 and sec(pi)/3.

OpenStudy (anonymous):

For the first one, cos2(pi)/3, would i simply input cos(120) and get the answer?

OpenStudy (kropot72):

Yes, if your calculator is set to degrees. If you calculator is set to radians you can enter cos((2 * pi)/3)

OpenStudy (anonymous):

I got -.5

OpenStudy (anonymous):

I think it needs to be in radical form

OpenStudy (anonymous):

which equals to -(pi)/6?

OpenStudy (anonymous):

oh thats wrong I meant radical 3/2

OpenStudy (anonymous):

Im not sure actually, please explain

OpenStudy (kropot72):

cos 120 = cos (180 - 60) = -cos 60 = -0.5

OpenStudy (anonymous):

How do I figure out -0.5 in radical form?

OpenStudy (kropot72):

\[-0.5=-\frac{1}{2}\] A radical form does not exist for 1/2

OpenStudy (anonymous):

Never mind, I figured it out, back to the other topic, if cot theta=0.3276, what is the radian measure? Is it 1.2542?

OpenStudy (kropot72):

You asked "how do I find the exact function value of an angle that is given in radians?" The simplest way is to set your calculator to radians. Then to find cos 2(pi)/3 you enter cos((2 * pi)/3) and should get the answer -0.5.

OpenStudy (anonymous):

Yea, thanx again, I actually noticed that I went back and read what you wrote, do you think you could address the last question I asked which can be found above.

OpenStudy (kropot72):

cot theta = 0.3276. Therefore tan theta = 1/0.3276. \[\theta=\tan^{-1} \frac{1}{0.3276}\] When you have found the value of theta in degrees you convert to radians by dividing by 180 and multiplying the result by pi.

OpenStudy (anonymous):

In that case, my answer is right.

OpenStudy (anonymous):

If f(x)=sin(1/3x), find f(pi/2)

OpenStudy (anonymous):

Is the answer 92?

OpenStudy (kropot72):

The answer is not correct. The maximum value of the sine function is 1, therefore 92 is not correct.

OpenStudy (anonymous):

Is it .9993?

OpenStudy (kropot72):

If f(x)=sin(1/3x), find f(pi/2) Is the following correct? \[f(x)=\sin \frac{1}{3}x,\ find\ f(\frac{\pi}{2})\]

OpenStudy (anonymous):

That is correct, the 1/3x is in parenthesis but that doesn't make a difference.

OpenStudy (kropot72):

Is the following correct? \[f(x)=\sin (\frac{1}{3x})\]

OpenStudy (anonymous):

the x is lined up with the division sign, but thats all, everything else is fine

OpenStudy (kropot72):

\[f(x)=\sin (\frac{1}{3}x)\] \[f(\frac{\pi}{2})=\sin (\frac{1}{3}\times \frac{\pi}{2})=\sin (\frac{\pi}{6})\] \[\frac{\pi}{6}\ radians=30\ degrees\] \[\sin 30degrees=you\ can\ calculate\]

OpenStudy (anonymous):

oh wow i got that initially but then changed it because I had thought that we had to have to same denominator but now I realize that's only when we are adding. Its easy thanks for showing me. I have one or two more questions i'd like to show you, it's sort of a new topic but I was absent on the day my teacher explained it to us, hopefully you can help me with it quikly. Thank you for everything, Im seriously grateful :)

OpenStudy (kropot72):

You're welcome :) Sorry, i must log out now. Please post as new questions.

OpenStudy (anonymous):

Theta is the radian measure of a central angle that intercepts an arc length, s, in a circle with a radius of length, r. If , s=6 and r=1, find theta

OpenStudy (anonymous):

@hartnn dude help me quikly please.

OpenStudy (anonymous):

@hartnn the question is the most recent reply

hartnn (hartnn):

s= r* theta where theta is in radians.

OpenStudy (anonymous):

They gave me s, does that mean theta is 6?

hartnn (hartnn):

yes, theta = 6 radians to be precise.

OpenStudy (anonymous):

the next one if if theta=4.5, and s=1.6, find r, soo would it be a negative number?

OpenStudy (anonymous):

1.6=r x 4.5

OpenStudy (anonymous):

how would i find r?

hartnn (hartnn):

length of arc s is just the product or radius r and theta angle in radians. so, s= r * theta 1.6 = r *4.5 [if theta =4.5 is given in radians] r = 1.6/4.5 =.... ?

OpenStudy (anonymous):

I got .355555555556

OpenStudy (anonymous):

is that what you got?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

Awsome just one, just one more, its if s=18, and theta=6(pi)/5, find r, ill try and do it, give me a sec

OpenStudy (anonymous):

.083333?

hartnn (hartnn):

how ?

OpenStudy (anonymous):

6(pi)/5=216

OpenStudy (anonymous):

and so i did 18/216

OpenStudy (anonymous):

and i got the answer

hartnn (hartnn):

don't convert angle in degrees! we need it in radians only to use that formula....

OpenStudy (anonymous):

Ohh so what do i do?

OpenStudy (anonymous):

Is it 4.7746...

hartnn (hartnn):

r= 18 / [6 pi /5] = [18*5]/ [6pi] yes! 4.7746 is correct! :)

OpenStudy (anonymous):

YES!!!!!! Lol thanks man! Seriously appreciate it!

hartnn (hartnn):

welcome ^_^